### v1

# Deconstructing What Makes a Good Optimizer for Language Models

Abstract
Training language models becomes increasingly expensive with scale, prompting numerous attempts to improve optimization efficiency. Despite these efforts, the Adam optimizer remains the most widely used, due to a prevailing view that it is the most effective approach. We aim to compare several optimization algorithms, including SGD, Adafactor, Adam, and Lion, in the context of autoregressive language modeling across a range of model sizes, hyperparameters, and architecture variants. Our findings indicate that, except for SGD, these algorithms all perform comparably both in their optimal performance and also in terms of how they fare across a wide range of hyperparameter choices. Our results suggest to practitioners that the choice of optimizer can be guided by practical considerations like memory constraints and ease of implementation, as no single algorithm emerged as a clear winner in terms of performance or stability to hyperparameter misspecification.
Given our findings, we further dissect these approaches, examining two simplified versions of Adam: a) signed momentum (Signum) which we see recovers both the performance and hyperparameter stability of Adam and b) Adalayer, a layerwise variant of Adam which we introduce to study Adam’s preconditioning. Examining Adalayer leads us to the conclusion that the largest impact of Adam’s preconditioning is restricted to the last layer and LayerNorm parameters, and, perhaps surprisingly, the remaining layers can be trained with SGD.
1 Introduction
As language model architectures increase in scale, pretraining becomes more expensive. In response, numerous efforts have been made to design efficient optimizers to mitigate these costs, and yet Adam [Kingma and Ba, 2015] remains the primary optimizer used for training language models. This persistent preference for Adam is rooted in an underlying belief that Adam generally outperforms alternative optimization algorithms. Although newly proposed optimizers run ablations to demonstrate superior performance to Adam for select architectures and tasks [Liu et al., 2024, Chen et al., 2023], there is no consensus among the literature about the relative performance of these optimizers. In fact, to the best of our knowledge, [Kaddour et al., 2024] is the only work comparing these optimizers but in the context of masked language modeling and at a single model scale.
In this work, we perform a comprehensive sweep for training autoregressive language models across different optimizers, hyperparameters, architectures, and scale. Along with looking at optimal performance, we argue that due to the difficulty of hyperparameter tuning with increasing scale [Yang et al., 2021], the stability of performance with respect to hyperparameter choices is equally important. Prior work has explored the learning rate stability of Adam [Wortsman et al., 2024]. We extend this investigation to include the stability of multiple optimizers with respect to various hyperparameter choices. Surprisingly, we find that multiple optimizers introduced in the literature after Adam—such as Lion Chen et al. [2023] and Adafactor (with momentum) Shazeer and Stern [2018], Zhai et al. [2022]—demonstrate robustness comparable to Adam and significantly superior to SGD. Figure 1 illustrates the remarkable similarity in performance and robustness of these optimizers across different learning rates and across multiple model scales (150m, 300m, and 600m parameters). This challenges the prevailing notion that Adam should be the default optimizer, where we see no single algorithm emerged as a clear winner in terms of performance or hyperparameter stability.
Following our initial ablations, we wish to identify the essential components of these optimizers that facilitate performance and stability. Thus, we conduct a series of investigations of simplified versions of these algorithms. We study signed momentum (Signum), a special case of Lion. Prior works have also studied its similarities to Adam [Balles and Hennig, 2018]. We find that Signum also recovers the stability and performance exhibited by Adam. This finding aligns with recent work [Kunstner et al., 2023], suggesting that the primary distinction between SGD and Adam is driven by Adam’s resemblance to signSGD.
To further understand the role of preconditioning of various network parameters, we study Adalayer, which performs preconditioning on a per-layer basis. We empirically demonstrate that this variant nearly recovers the stability and performance of the other optimizers in previous ablations. Through empirical studies of Adalayer and its variants, we show that while adapting the parameters of the last layer and LayerNorm parameters in a transformer is necessary to achieve stability and performance, we can actually train the remaining parameters (a vast majority of the network) with SGD.
Main contributions
We empirically study the stability to hyperparameters of various optimization algorithms including SGD, Adam, Lion and Adafactor, showing that with the exception of SGD, these optimizers are comparable in terms of both performance and hyperparameter stability. This holds across multiple scales (150m, 300m, 600m) and across two transformer architecture variants (Section 3).
To dive into the reasons behind the stability and performance of these algorithms, we empirically examine signed momentum (Signum), and show that it recovers their stability and performance (Section 3.6).
We study a coarser variant of Adam called Adalayer, that does per-layer preconditioning and recovers much of the stability and performance exhibited by Adam (Section 4.1).
Through an empirical study of Adalayer and its variants, we establish that adaptivity is only necessary for the last layer and LayerNorm parameters, while the remaining parameters can be trained with SGD (Section 4.2).
2 Related work
One closely related work to ours is Wortsman et al. [2024], which explores the stability of Adam with respect to learning rate. We extend the comparison to other optimizers including SGD, Lion and Adafactor, as well as other hyperparameters including momentum and weight decay.
Optimizers: SGD [Robbins and Monro, 1951] had been the workhorse optimizer for deep learning until 2015, when Adam [Kingma and Ba, 2015] was introduced. Adam is a diagonal preconditioning algorithm that maintains a per-parameter learning rate. Over time, coarser variants of Adam have been proposed, which do not explicitly maintain a learning rate per parameter. Adafactor [Shazeer and Stern, 2018, Zhai et al., 2022] maintains a rank-1 approximation of the preconditioner matrix of Adam. Prior works have also explored the similarity of Adam with variants of Signum [Balles and Hennig, 2018], and recently, a close variant of Signum, called Lion [Chen et al., 2023], was discovered using symbolic search over algorithms. Some other optimizers that have recently gained increasing attention from the community include Shampoo Gupta et al. [2018] and Sophia [Liu et al., 2024].
Adam and Signum: Many works have explored the relationship between Adam and variants of Signum [Balles and Hennig, 2018, Balles et al., 2020, Kunstner et al., 2023] and empirically demonstrated that Signum (or its close variants) generally performs comparably to Adam. [Balles et al., 2020] also argued that signSGD generally performs better when the Hessian is close to diagonal, however, it is unclear if this holds for practical settings. Kunstner et al. [2023] recently demonstrated that Adam and a close variant of Signum exhibit similar performance on a variety of datasets including WikiText-2 [Merity et al., 2017] and SQuAD [Rajpurkar et al., 2016]. However, in contrast with our work, all of these are restricted to the setting of vision or masked language modeling, and generally do not sweep over multiple hyperparameters.
Layerwise or blockwise Adam: We study Adalayer, a layerwise version of Adam. This is a special case of the BAGM optimizer Zheng and Kwok [2019], specifically BAGM B.1. Similar algorithms have also been studied by [Ginsburg et al., 2019, Agarwal et al., 2020, Liu et al., 2021, Zhang et al., 2024b]. In particular, concurrent to our work, Zhang et al. [2024b] propose an algorithm termed Adam-mini, which closely tracks a modified version of Adalayer (called Adalayer*), and demonstrate comparable performance to AdamW. Note that, in our work, Adalayer* is introduced to understand the role played by preconditioning in Adam, and we do not specifically focus on the final performance. [Zhang et al., 2024a] empirically study the Hessian spectrum of transformers at initialization and find it to be more heterogeneous across layers as compared to ResNets. They argue that this heterogeneity is evidence towards the importance of Adam in training transformers. In contrast our results (Section 4.2) show that all layers except last layer and LayerNorm layers can be trained with SGD.
Other related works: For vision transformers, in the fine-tuning phase, Kumar et al. [2024] show that using SGD with frozen embedding parameters leads to competitive performance with Adam. Jelassi et al. [2022] explore the similarity between Adam and normalized gradient descent [Nesterov, 2004] and show that normalized gradient descent on GANs does not suffer from mode collapse, while SGD does. Jiang et al. [2023] empirically demonstrate that Adam steers the parameter trajectory towards better-conditioned regions than SGD. Pan and Li [2022] also show that the parameter trajectory of Adam exhibits much higher directional smoothness than that of SGD. Ahn et al. [2024] show that the performance gap between Adam and SGD exacerbates with depth of the network.
3 Comparing Optimizers Across Hyperparameters, Architectures and Scale
3.1 Methodology
To conduct our experiments, we start with hyperparameters recommended by previous work (e.g., ). We initially perform a learning rate sweep to identify the optimal learning rate. After determining the optimal learning rate for each algorithm, we conduct one-dimensional sweeps for each of the other hyperparameters.
A limitation of this methodology is the potential neglect of "2D" interactions between hyperparameters. This is an important direction for future work, but beyond the computational budget of this project. For example, some parameters like batch size and learning rate indeed are likely to exhibits 2D interactions Shallue et al. [2019], Porian et al. [2024]. However, we argue that the 1D sweeps provide a tractable methodology that gives us useful signal about the hyperparameter stability of a variety of algorithms around the parameters that are common in practice.
3.2 Setup
We train language models on C4 tokenized with the T5 tokenizer [Raffel et al., 2020] and report results in terms of validation loss. As we discussed in the introduction, we argue that it is best to evaluate algorithms both in terms of the loss achieved by the best hyperparameters (performance) as well as the robustness across values of the hyperparameters (stability).
In this section we first present the results of sweeps across the two most sensitive hyperparameters: learning rate and momentum (). We sweep across five algorithms: Adam, Adafactor, Lion, Signum, and SGD. Further ablations of weight decay, warmup, , and for the 150m standard model can be found in Section 3.5. Before diving into the results, we provide some more details about the setup.
Algorithms.
We use the standard Pytorch implementation of AdamW [Paszke et al., 2019], the timm implementation of SGDW [Wightman, 2019], and the OLMo implementation of Lion [Groeneveld et al., 2024]. Following Zhai et al. [2022] we implement ourselves a modified version of Adafactor which maintains the factored estimates of second moments but has momentum i.e. it is equivalent to Adam with factored second moment estimates. Since Signum is equivalent to Lion with we reuse the OLMo implementation of Lion [Groeneveld et al., 2024] for it. We conducted experiments with the Sophia optimizer Liu et al. [2024] in Appendix C. However, since it does not outperform Signum (which can be achieved by setting in Sophia), we did not include it in other plots.
Models.
We start from the OLMo codebase [Groeneveld et al., 2024] and train decoder-only transformer models of three sizes: 150m, 300m, and 600m, where the parameter count refers to non-embedding parameters. The models have widths of 1024, 1024, and 1408 and depths of 12, 24, 24. The MLP hidden dimension is 4x of the width. The activation function is GeLU [Hendrycks and Gimpel, 2016]. We use RoPE positional encodings [Su et al., 2024]. Attention heads are always dimension 64. We use PyTorch default LayerNorm. Following Wortsman et al. [2024] we do not learn biases for the linear layers or LayerNorms. We train in mixed precision with bfloat16.
Training variants.
We note that Wortsman et al. [2024] observe that QK LayerNorm [Dehghani et al., 2023] and z-loss [Chowdhery et al., 2023] can have substantial effects on the stability of model training. As such, we consider two variants in our experiments: standard which refers to a model with QK LayerNorms and z-loss with coefficient 1e-4, and no QK norm or z-loss which refers to the same model without the QK norm layers or the z-loss.
Token counts.
For all models, we use a batch size of 256 and sequence length of 512 (as in Wortsman et al. [2024]). We default to training models for the approximately “chinchilla optimal” number of tokens that is 20 times the number of parameters. Explicitly, this means for the 150m models we train for 25k steps or 3.3b tokens. The 300m models are trained for 50k steps, the 600m models are trained for 100k steps and the 150m-long models are also trained for 100k steps.
Other hyperparameters.
We default to using 0 weight decay. We default to using a learning rate schedule with 10% of the training steps for warmup and then cosine decay with a minimum that is 10% of the maximum learning rate. We default to and following Wortsman et al. [2024]. These parameters are ablated in Section 3.5.
3.3 Sweeping learning rates
First, we sweep over the most important hyperparameter: learning rate. Note, in all of these sweeps over learning rate we set for all algorithms except for SGD, where we set . As we will see in the following subsection, SGD is more sensitive to the momentum hyperparameters and requires more momentum to be competitive with the other optimizers.
Main results for our standard architecture across three scales are presented in Footnote 1. Note that the x-axis shifts the learning rates to align the optimal learning rates across algorithms. In terms of absolute learning rates, we sweep in multiples of from 1e-4 to 1 for Adam and Adafactor, from 1e-5 to 1e-1 for Lion and Signum, and from 1e-3 to 10 for SGD.
The key takeaway is that not only do the algorithms achieve similar performance at the optimal learning rate, but the learning rate stability itself is similar across algorithms and scales. The one exception is SGD, which is worse both in terms of optimal performance and in terms of stability.
Further ablations are presented in Figure 2 and Figure 3 illustrating performance for models with no QK norm or z-loss and 4x longer training time respectively. While we find that the architecture choices can clearly impact the amount of stability to learning rate, the cross-algorithm comparisons remain the same: Adafactor and Lion are competitive with Adam, while SGD is worse both in terms of performance and stability to learning rate. Similarly, training for longer can improve performance and stability to learning rate, but does not change the high-level cross-algorithm comparisons.
Takeaway:
performance and stability to learning rate are comparable across the non-SGD algorithms that we tested.
3.4 Sweeping momentum
Now we also sweep across momentum values (i.e. )222Note that in Lion, both and can be thought of as different types of “momentum” with being the “one-step” momentum and the “long-term” momentum. For consistency, we only sweep here and provide further ablations in the Appendix.. To do this sweep we fix the per-algorithm learning rate to be the optimal learning rate from the corresponding learning rate sweep.
Results are presented in Figure 4. We observe that across various settings, the robustness to is similar across the non-SGD algorithms when we stay in the range of momentums between 0.8 and 0.98. However, for high Lion is better and low Adam and Adafactor are better. Again we observe SGD being very sensitive to momentum.
Takeaway:
performance and stability to momentum are comparable across the non-SGD algorithms that we tested if we stay within the usual range of momentum values.
3.5 Additional hyperparameter sweeps
We also sweep over a variety of other hyperparameters in Figure 5 using the best per-algorithm learning rate and momentum. We observe that SGD is less stable with respect to weight decay and warmup length. And while it is possible to get small benefits from higher weight decay, longer warmup, and higher than our defaults, the algorithms are much more stable to these parameters than learning rate and momentum.
Takeaway:
generally algorithms are more stable with respect to other hyperparameters and the possible gains in performance are relatively small compared to learning rate and momentum.
3.6 Signum recovers the performance and stability of Adam
In Figure 1 we observe that Adam and Signum have similar performance and stability for language modeling, even at scale. The following lemma from prior work[Balles and Hennig, 2018] shows that Adam performs variance-adjusted sign gradient descent.
Lemma 1 ( Balles and Hennig [2018]).
Consider a parameter with a history of gradients . Let be the random variable that is equal to with probability and be the random variable that is equal to with probability . The Adam update and the Signum update are related by
If then in Lemma 1 and hence the ratio of Adam and Signum updates is equal to the ratio of the mean and the square root of second moment of . Intuitively, this holds because when , first moment estimates of Signum and Adam, and second moment estimates of Adam, average the previous gradients with same coefficients (). This intuitively suggests that when , Adam and Signum may behave similarly. This motivates the conjecture that the main benefit of Adam over Signum is the fact that in Adam, can be varied independently of . In Figure 1 we have and which are close, and as pointed out earlier, both optimizers have similar performance and stability.
We examine this hypothesis further in Figure 6 by varying and setting , and again find that Signum and Adam behave very similarly. However, we also note that when we vary for Adam while fixing we get more stability for as compared to Signum.
Takeaway: With Adam and Signum behave similarly and the standard setting for training language models () is close to this.
4 Investigating the key factors for optimizer stability and performance
Ablations in the previous section revealed the striking similarity in performance and stability across multiple optimizers compared to Adam. In this section we seek a finer-grained understanding of the role of preconditioning for adaptive optimizers in achieving both performance and stability. Adam and its other variants are designed to have a high degree of adaptivity at a fine granularity (per-parameter learning rates) throughout the training process. This adaptivity is often credited with the stability and robust performance observed in these optimizers. However, a critical question arises: to what extent is this adaptivity needed for different parameters of the network? By dissecting the fundamental components that ensure both performance and stability, we aim to discern whether simpler optimizers like SGD can achieve similar benefits with minimal modifications. Since higher momentum can often play the same role as a better preconditioner and to have all algorithms on an equal footing, we will fix for all optimizers in this section.
The main optimizer we study in this section is a “layer-wise” variant of Adam, which we coin as ‘Adalayer’. We use Adalayer for our investigations because it lends a greater ease of understanding compared to full-fledged Adam in identifying parts of the network which may be particularly critical for optimizer performance and stability. Note that this layerwise variant is a special case of a previously known optimizer called Blockwise Adaptive Gradient with Momentum (BAGM) [Zheng and Kwok, 2019].
4.1 Adalayer
To study the behavior of adaptive optimizers like Adam, we begin with describing a layer-wise version of Adam which we refer to as Adalayer. Adam, Adafactor and Adalayer all (approximately) store the diagonal second moment matrix, but with coarser and coarser granularity; for a layer of dimension , Adam explicitly maintains the second moment matrix using parameters in the shape of a matrix. Adafactor stores row and column averages of the second moment matrix which serve as a rank-1 approximation to the second moment matrix. Finally, Adalayer stores a single scalar which is the average of the second moment matrix. We will later consider a generalization of Adalayer where instead of averaging second moment over a layer we will average it over a “block” of parameters which can be a subset of a layer. We note that similar algorithms have been studied before by Ginsburg et al. [2019], Agarwal et al. [2020] but we choose to study this variant since it is a direct analogue of Adam and Adafactor. A simplified version of Adalayer optimizer is given in Algorithm 1, other details such as bias correction are kept same as that for Adam.
In Figure 7 we study the behavior of Adalayer across learning rates. To preserve the correspondence with Adam we fix other hyperparameters to be the same: , and . We find that Adalayer has better performance than SGD333Note that in this section we are setting for all optimizers., but it performs worse than Adam and also lacks Adam’s stability across learning rates. The major difference between Adam and Adalayer is the preconditioning done by Adam within a layer. Intuitively, this preconditioning will have large effects in layers where we expect different weights within a layer to have different gradient scales. The first candidate for such a layer is the last layer, since different tokens have widely different frequencies leading to different gradient scales. To test this hypothesis, we run a corrected version444We note that this reasoning also applies to the first layer, but in our ablations applying this correction the first layer did not make a significant difference. of Adalayer where we treat the set of weights feeding into a logit as a separate block. We henceforth refer to Adalayer with this correction as Adalayer*. This is plotted in Figure 7 and we observe that Adalayer* almost recovers the performance as well as a large fraction of the stability of Adam.
To study how Adalayer* preconditions the network, we plot effective learning rates used for different logits by Adalayer* in Figure 8 (Left). Here, the effective learning rate for a layer in the network is . We find that Adalayer* indeed uses vastly different learning rates for different logits, supporting our hypothesis that preconditioning weight in different logits separately is important for performance and stability.
4.2 Most language model parameters can be trained with SGD
The results using Adalayer* in the previous section suggest that all layers except the last layer only need a iteration-dependent scalar correction to their learning rate. We now ask a stronger question: do we need these scales at all? Or can we train the remaining layers with SGD? This hypothesis is supported by looking at Figure 8 (middle) where we observe that the learning rates for different matrix layers (except the last layer) assigned by Adalayer* are remarkably similar.
To test this, we train the last layer with Adalayer* (fixing a learning rate of ) and the rest of the layers with SGD, both with . In Figure 9 (Left) we show the results while sweeping over SGD learning rates from to . While this improves upon the performance of SGD, we do not recover stability of the Adalayer* and Adam baselines. We trace this instability to LayerNorm blocks: Figure 8 (Left) shows that the effective learning rates for the LayerNorm blocks are much smaller, which suggests that they may destabilize at higher SGD learning rates. To ameliorate this, in Figure 9 (Middle and Right) we add LayerNorm parameters to those being trained with Adalayer* and find that this is sufficient to recover both performance and stability of Adalayer*. For the larger 300m model, we find that this even exceeds the performance of Adalayer*.
In the Appendix, we conduct further experiments investigating these ‘hybrid’ variants of SGD. Another plausible remedy to address the small effective learning rates of the LayerNorm blocks is to simply not train these LayerNorm parameters. In Appendix A, Figure 11, we show that this improves performance and stability relative to SGD but does not recover Adalayer* performance. Furthermore, in Figure 12 we train 150m and 300m models using SGD on the matrix blocks and AdaFactor on the last layer and LayerNorm blocks; we find that performance and stability is comparable, but does not exceed that of Adafactor on the entire network.
Note that a caveat of the above results is that we have introduced an additional hyperparameter— SGD learning rate, which we are sweeping over— while keeping the Adalayer* learning rate in the last layer and layer norm layers fixed. While decoupling the learning rates here is needed (due to SGD’s performant learning rates being orders of magnitude higher than that of Adalayer*), this may be responsible for the observed stability. To address this, we perform the following experiment: we train all the layers with Adalayer* (sweeping over learning rate ) but we stop updating the second moment estimates for all layers except the last layer and LayerNorm blocks after initialization. This implies that these layers are effectively being trained by SGD with a fixed learning rate, though unlike the above results, these learning rates are different for different layers. We implement this by passing 1000 batches to initialized 150m and 300m models to obtain second moment estimates for all layers without letting the model take a gradient step, and then allowing the model to train as normal under the same settings as all of our ablations. As in our previous investigation, we fix other hyperparameters to be the same: , and .
In Figure 10 we show the resulting learning rate sweep for freezing Adalayer* learning rate scales at initialization (with the exception of the last layer and LayerNorm). Surprisingly, we find for the 150m model that we can almost entirely recover the stability and performance of Adalayer*. For the 300m model we also match or exceed the performance of Adalayer*, and even nearly match the peak performance of Adam. Note again that this sweeps learning rate across all network parameters. This provides further evidence for the importance of adaptivity in the last layer and LayerNorm, where in contrast we could used fixed ratios from initialization for all other parameters to recover the performance and stability of Adalayer*. In Appendix B we present additional experiments which include allowing Adalayer* to update for only one of the last layer or LayerNorm, or turning off LayerNorm training. None of these partial modifications fully recovers performance and stability.
5 Discussion and Limitations
After a comprehensive comparison of a variety of optimizers for language modeling, we have found that many optimizers seem to be roughly equivalent both in terms of optimal performance and hyperparameter stability. Diving deeper, we have shown that most of the benefits of adaptive optimizers arise from their treatment of the last layer and LayerNorm parameters. Of course, there are several limitations to our study including the fact that due to computational constraints we only ablate a few architecture decisions, that we only consider one dimensional hyperparameter sweeps, we fix batch size, and that we limit our study to language modeling with a single dataset. Despite these limitations, we believe that the study sheds new light on the fundamentals of optimization for language modeling and suggests that optimizer choice may not be the optimal point of intervention for increasing efficiency.
Acknowledgments
SK, DM and RZ acknowledges support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. NV, DM and RZ are supported by a Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199.
Appendix A Additional experiments: SGD + adaptive variants (Adalayer*, Adafactor)
In this section, we report additional experiments involving training language models with SGD on a fraction of the models’ parameters and an adaptive optimizer on the remaining parameters. Firstly, given that the effective learning rates of the LayerNorm blocks were observed to be small in Figure 8, it is reasonable to ask whether training the LayerNorm parameters is necessary at all; in Figure 11, we show results for training 150m and 300m models using Adalayer* only on the last layer, using SGD on all other matrix blocks, and turning off training for the LayerNorm parameters. This indeed yields greater stability in comparison to Figure 9 (Left) but does not fully recover the performance of Adalayer*, indicating that training LayerNorm parameters helps with performance, which seems more pronounced in the larger model.
We saw in Figure 9 (Middle, Right) that using Adalayer* on only the last layer and LayerNorm parameters sufficed to recover or exceed the performance of Adalayer*. In Figure 12, we report a learning rate sweep over the analogous experiment but using Adafactor on the last layer and LayerNorm parameters with a fixed learning rate. For the 150m model, using a learning rate of with Adafactor yielded better performance than Adafactor for low learning rates, and is comparable in terms of performance and stability. For the 300m model, the difference between Adafactor and our ‘hybrid’ optimizer is more distinct at higher learning rates for fixed Adafactor learning rate and , but is comparable until the peak validation loss.
Appendix B Additional experiments: freezing Adalayer learning rate ratios
In this section, we report additional experiments exploring whether the results involving frozen Adalayer* in Section 4.2 need both last layer and LayerNorm adaptivity. We show that this is indeed the case by conducting the same sweep for frozen Adalayer* while trying to also freeze the learning rate ratios for last layer or LayerNorm parameters as well. In Figure 13, we show that fixing initialized learning rate ratios for all layers does not reach peak performance of Adalayer*, nor does it exhibit stability. In Figure 14, we show that either continuing to update the LayerNorm parameters or the last layer parameters can achieve the peak performance of Adalayer* but is still unstable. Finally, we show results for turning off LayerNorm training while fixing learning rate ratios (with the exception of the last layer) in Figure 15. We conclude that it is necessary to maintain adaptivity for both the last layer and LayerNorm parameters, but understanding why the fixed ratios do not suffice would be an interesting question for future work.
Appendix C Sophia
In this section, we compare Sophia [Liu et al., 2024] to Signum. Note that Signum is a special case of Sophia, achieved by setting . We find that Sophia does not outperform Signum. No significant change in performance was observed when transferring the hyperparameters suggested by Liu et al. [2024] (eg. , , , weight decay), nor when additionally scaling attention by the inverse of layer index which was used in the original Sophia implementation.